Copper exists in face centered cubic lattice with unit cell edge of 361pm . What is the size of the largest atom , which would fit into ocathedral voids without disturbimg the lattice

Grade:12

2 Answers

Sunil Kumar FP
9 years ago
As the copper crystallises in FCC.therefore
edgelength=a=361pm
The radius of copper atom is 4RCu=sqrt2 a=Rcu=sqrt2 a/4=108.3 pm
here the largest size of atom that can be fitted into octahedral void is
2(Rcu +r-)-=a
r-=180.5-108.3 pm=72.2 pm

Sunil Kumar FP
9 years ago
Copper exists in face centered cubic lattice with unit cell edge of 361pm . What is the size of the largest atom , which would fit into ocathedral voids without disturbimg the lattice

in the above the radius of Cu=a*sqrt2/4=361*1.414/4=127 pm
therrefore size of largest atom=180.5-127 pm=52.8 pm